The formation of CsCl from Cs(s) and Cl2(g) involves the following steps:

Cs(s)?Cs(g)

12Cl2(g)?Cl(g)

Cs(g)?Cs+(g)+e?

Cl(g)+e??Cl?(g)

Cs+(g)+Cl?(g)?CsCl(s)

Which of these steps absorb energy and which release energy?
Answer #1
Concepts and reason

The energy required to break ionic solid into its constituent gaseous ions is termed as lattice energy. And the stability of ionic solids depends upon the lattice energy.

This lattice energy is calculated by adding or subtracting different energies of the steps involved during the formation of ionic compound.

Fundamentals

Energy is the property of a substance. It can only be transferred from one substance to another but cannot be destroyed.

If a substance absorbs energy, it is termed as endothermic energy. And if a substance releases energy, it is termed as exothermic energy.

The reaction is

Cs(s)Cs(g){\rm{Cs}}\left( s \right) \to {\rm{Cs}}\left( g \right)

Here, solid Cs is converted into gaseous Cs, therefore, energy is absorbed.

The reaction is

Cl2(g)2Cl(g){\rm{C}}{{\rm{l}}_2}\left( g \right) \to 2{\rm{Cl}}\left( g \right)

Here, the bond between two Cl atoms is breaking which needs energy, therefore, energy is absorbed.

The reaction is

Cs(g)Cs+(g)+1e{\rm{Cs}}\left( g \right) \to {\rm{C}}{{\rm{s}}^ + }\left( g \right) + 1{e^ - }

Here, ionization is taking place which is an endothermic reaction.

The reaction is

Cl(g)+1eCl(g){\rm{Cl}}\left( g \right) + 1{e^ - } \to {\rm{C}}{{\rm{l}}^ - }\left( g \right)

Here, energy is released as an electron is added into an atom.

The reaction is

Cs+(g)+Cl(g)CsCl(s){\rm{C}}{{\rm{s}}^ + }\left( g \right) + {\rm{C}}{{\rm{l}}^ - }\left( g \right) \to {\rm{CsCl}}\left( s \right)

Here, energy is released as oppositely charged ions experiences attractive forces due to the formation of ionic solid.

Ans:

The reaction Cs(s)Cs(g){\rm{Cs}}\left( s \right) \to {\rm{Cs}}\left( g \right) is endothermic, that is, energy is absorbed.

The reaction Cl2(g)2Cl(g){\rm{C}}{{\rm{l}}_2}\left( g \right) \to 2{\rm{Cl}}\left( g \right) is endothermic, that is, energy is absorbed.

The reaction Cs(g)Cs+(g)+1e{\rm{Cs}}\left( g \right) \to {\rm{C}}{{\rm{s}}^ + }\left( g \right) + 1{e^ - } is endothermic, that is, energy is absorbed.

The reaction Cl(g)+1eCl(g){\rm{Cl}}\left( g \right) + 1{e^ - } \to {\rm{C}}{{\rm{l}}^ - }\left( g \right) is exothermic, that is, energy is released.

The reaction Cs+(g)+Cl(g)CsCl(s){\rm{C}}{{\rm{s}}^ + }\left( g \right) + {\rm{C}}{{\rm{l}}^ - }\left( g \right) \to {\rm{CsCl}}\left( s \right) is exothermic, that is, energy is released.